OpenStudy (anonymous):

Calculate 40^13 (mod85)

Nnesha (nnesha):

@ganeshie8 "mod" <----- :D

ganeshie8 (ganeshie8):

lol

OpenStudy (inkyvoyd):

mm so you guys are covering modular exponentiation?

OpenStudy (inkyvoyd):

there are two key modular arithmetic laws that will help you here, namely ab mod c is congruent to a mod c * b mod c

OpenStudy (inkyvoyd):

or maybe that was the only law lol

Nnesha (nnesha):

^_^ idk i just saw him doing "mod" problems so i just tagged .....

OpenStudy (inkyvoyd):

@danyboy9169 are you there?

OpenStudy (anonymous):

yes, just few hrs old on Open Study so trying to understand how this works

OpenStudy (mathmath333):

its easy let me post a solution

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{40^{13} \pmod {85}\hspace{.33em}\\~\\ =40\cdot 40^{12} \pmod {5\times 17}\hspace{.33em}\\~\\ =8\cdot 40^{12} \pmod { 17}\hspace{.33em}\\~\\ =8\cdot (17\times 2+6)^{12} \pmod { 17}\hspace{.33em}\\~\\ =8\cdot (6)^{12} \pmod { 17}\hspace{.33em}\\~\\ =8\cdot (6^2)^{6} \pmod { 17}\hspace{.33em}\\~\\ =8\cdot (36)^{6} \pmod { 17}\hspace{.33em}\\~\\ =8\cdot (17\times 2+2)^{6} \pmod { 17}\hspace{.33em}\\~\\ =8\cdot (2)^{6} \pmod { 17}\hspace{.33em}\\~\\ =8\cdot (2^2)^{3} \pmod { 17}\hspace{.33em}\\~\\ =8\cdot (4)^{3} \pmod { 17}\hspace{.33em}\\~\\ =8\cdot4\cdot (4)^{2} \pmod { 17}\hspace{.33em}\\~\\ =8\cdot4\cdot 16 \pmod { 17}\hspace{.33em}\\~\\ =8\cdot4\cdot (17-1) \pmod { 17}\hspace{.33em}\\~\\ =8\cdot4\cdot (-1) \pmod { 17}\hspace{.33em}\\~\\ =32\cdot (-1) \pmod { 17}\hspace{.33em}\\~\\ =(17\times 2-2)\cdot (-1) \pmod { 17}\hspace{.33em}\\~\\ =(-2)\cdot (-1) \pmod { 17}\hspace{.33em}\\~\\ =2 \pmod { 17}\hspace{.33em}\\~\\ \normalsize \text{we have calculated mod 17 but the question asks mod 85 , so} \hspace{.33em}\\~\\ \normalsize \text{but 17*5=85} \hspace{.33em}\\~\\ =2\times 5 \pmod { 17\times 5}\hspace{.33em}\\~\\ =10 \pmod { 85}\hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

ask for doubt if any in a step or 2

OpenStudy (thomas5267):

TBH I rather multiply 40 then reduce it and do it 13 times than do it your way. The steps required are too complex for me.

OpenStudy (thomas5267):

I reread it and it is much better than when I first read it lol.

ganeshie8 (ganeshie8):

\[8\times 40^{12}\equiv 2 \pmod{17} \implies 5\times 8\times 40^{12} \equiv 5\times 2 \pmod{5\times 17} \]

OpenStudy (thomas5267):

Something seems wrong though. I don't think \(2\pmod {17}=10\pmod {85}\). For example, \(19\equiv2\pmod {17}\) but clearly \(19\neq10\pmod{85}\)

ganeshie8 (ganeshie8):

x/5 = 2 (mod 17) multiply 5 through out x = 10 (mod 85)

ganeshie8 (ganeshie8):

\[2 \pmod{17} \ne 10 \pmod{85}\]

ganeshie8 (ganeshie8):

you're right abot that but i dont see anyone saying they are eqal in this thread ?

OpenStudy (thomas5267):

The last five lines of mathmath333 proof.

ganeshie8 (ganeshie8):

I see... he meant this i think : we want to find the remainder \(x\equiv 40^{13} \pmod{85}\) divide through out by 5 \(x/5\equiv 8\cdot 40^{12} \pmod{17}\) \(\cdots\) \(x/5\equiv 2 \pmod{17}\) multiply by 5 through out \(x\equiv 10 \pmod{85}\)

ganeshie8 (ganeshie8):

he just divided 5 in the start so that he can work easily in a smaller moduli in the end he miltiplied everything by 5 to keep things same looks good to me :)

OpenStudy (thomas5267):

If \(x\equiv a\pmod n\), does \(bx\equiv ba \pmod {bn}\)?

ganeshie8 (ganeshie8):

yes lets prove it

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{40^{13} \pmod {85}\hspace{.33em}\\~\\ =40\cdot 40^{12} \pmod {\color{red}{5}\times 17}\hspace{.33em}\\~\\ =8\cdot 40^{12} \pmod { 17}\hspace{.33em}\\~\\ =8\cdot (17\times 2+6)^{12} \pmod { 17}\hspace{.33em}\\~\\ =8\cdot (6)^{12} \pmod { 17}\hspace{.33em}\\~\\ =8\cdot (6^2)^{6} \pmod { 17}\hspace{.33em}\\~\\ =8\cdot (36)^{6} \pmod { 17}\hspace{.33em}\\~\\ =8\cdot (17\times 2+2)^{6} \pmod { 17}\hspace{.33em}\\~\\ =8\cdot (2)^{6} \pmod { 17}\hspace{.33em}\\~\\ =8\cdot (2^2)^{3} \pmod { 17}\hspace{.33em}\\~\\ =8\cdot (4)^{3} \pmod { 17}\hspace{.33em}\\~\\ =8\cdot4\cdot (4)^{2} \pmod { 17}\hspace{.33em}\\~\\ =8\cdot4\cdot 16 \pmod { 17}\hspace{.33em}\\~\\ =8\cdot4\cdot (17-1) \pmod { 17}\hspace{.33em}\\~\\ =8\cdot4\cdot (-1) \pmod { 17}\hspace{.33em}\\~\\ =32\cdot (-1) \pmod { 17}\hspace{.33em}\\~\\ =(17\times 2-2)\cdot (-1) \pmod { 17}\hspace{.33em}\\~\\ =(-2)\cdot (-1) \pmod { 17}\hspace{.33em}\\~\\ =2 \pmod { 17}\hspace{.33em}\\~\\ \normalsize \text{we have calculated mod 17 but the question asks mod 85 , so} \hspace{.33em}\\~\\ \normalsize \text{but 17*5=85} \hspace{.33em}\\~\\ =2\times 5 \pmod { 17\times \color{red}{5}}\hspace{.33em}\\~\\ =10 \pmod { 85}\hspace{.33em}\\~\\ }\end{align}\)

ganeshie8 (ganeshie8):

\(x\equiv a \pmod{n} \implies n|(x-a) \implies bn|b(x-a)~~ \implies b(x-a) \equiv 0 \pmod{bn}\)

OpenStudy (mathmath333):

first i removed 5 and and then gave it back to make life easy

OpenStudy (mathmath333):

can also use the chineese remainder theorem here but that is not need since i can cancel 5

ganeshie8 (ganeshie8):

Exactly !

OpenStudy (mathmath333):

proof from wolfram http://www.wolframalpha.com/input/?i=40%5E%2813%29+mod+85%3D

OpenStudy (mathmath333):

the actual calculation is very short , but i extended it step by step to understand comprehensively

OpenStudy (mathmath333):

\(19\equiv 2 \pmod {17}\) and \(19*5=95\equiv 2\times 5\equiv 10 \pmod {85}\)

OpenStudy (anonymous):

Wow!! you guys are amazing.. I am new to this study group, am amazed the responses I have had. gr!8. - thanks We can use the rules of exponents to write 40^13 (mod 85) as (40^2)^6 (mod 85. ) Since 40^2 = 1600 = equiv 70 (mod 85), we can write our expression as 70^6 .40(mod 85) Applying rules of exponents again, we have (70^2)^3* 40 mod 85. With 70^2 = 4900 = equiv 55(mod 85), our expression is now 55^3 * 40 (mod 85) We can also rewrite this as 55^2 * 55 * 40 (mod 85) With 55^2 = 3025 =equiv 50 (mod 85) and 55 * 40 = 2200 =equiv 75 (mod 85), we can simplify the expression again to 50 * 75 (mod 85) We calculate 50* 75 = 3750=equiv 10 (mod 85.) Ans: 10

OpenStudy (mathmath333):

ur applying \(\mod 85 \) which is also good, but \(\mod 17 \) is definitely easy that \(\mod 85\)

OpenStudy (anonymous):

@mathmath333 I agree and thank all guys!